# ThinkSound: Chain-of-Thought Reasoning in Multimodal Large Language Models for Audio Generation and Editing

### v2

@@ Abstract
While end-to-end video-to-audio generation has greatly improved, producing high-fidelity audio that authentically captures the nuances of visual content remains challenging. Like professionals in the creative industries, such generation requires sophisticated reasoning about items such as visual dynamics, acoustic environments, and temporal relationships. We present ThinkSound, a novel framework that leverages Chain-of-Thought (CoT) reasoning to enable stepwise, interactive audio generation and editing for videos. Our approach decomposes the process into three complementary stages: foundational foley generation that creates semantically coherent soundscapes, interactive object-centric refinement through precise user interactions, and targeted editing guided by natural language instructions. At each stage, a multimodal large language model generates contextually aligned CoT reasoning that guides a unified audio foundation model. Furthermore, we introduce AudioCoT, a comprehensive dataset with structured reasoning annotations that establishes connections between visual content, textual descriptions, and sound synthesis. Experiments demonstrate that ThinkSound achieves state-of-the-art performance in video-to-audio generation across both audio metrics and CoT metrics and excels in out-of-distribution Movie Gen Audio benchmark. The project page is available at https://ThinkSound-Project.github.io.
@@ 1 Introduction
Figure 1: ThinkSound with CoT: (1) CoT-driven foley synthesis capturing semantic and temporal details, (2) interactive object-centric refinement for user control, and (3) targeted audio editing.
Generating realistic sound for video demands more than recognizing objects; it requires reasoning about complex visual dynamics and context – determining when an owl is chirping versus flapping its wings, identifying the subtle sway of tree branches, and synchronizing multiple sound events within a scene, as illustrated in Figure 1. Current end-to-end video-to-audio (V2A) generation systems (Luo et al., 2024; Xing et al., 2024; Zhang et al., 2024), while having largely improved, often struggle with this compositional complexity and contextual nuance. They may produce generic sounds or fail to synchronize precisely with subtle visual cues, thus limiting fidelity and user control.
Recent breakthroughs in Multimodal Large Language Models (MLLMs) (Lu et al., 2024; Liu et al., 2023b; Wang et al., 2024b) offer powerful capabilities in understanding and reasoning across multiple modalities. Concurrently, Chain-of-Thought (CoT) prompting (Wei et al., 2022; Zhang et al., 2023; Wang et al., 2025) has proven effective at eliciting structured, step-by-step reasoning from large models. These advancements create great potential to fundamentally rethink V2A by decomposing complex sound design into explicit reasoning steps and actionable synthesis instructions. The necessity for CoT in V2A becomes evident when examining the process of sound designers. They employ a multi-stage approach: analyzing visual content, then reasoning about acoustic properties, and finally synthesizing and refining sounds. End-to-end approaches compress the process into a single black-box transformation, losing the nuanced reasoning required for audio generation.
Some approaches have attempted to adopt MLLMs for multi-stage V2A. SonicVisionLM (Xie et al., 2024) converts video to textual captions using MLLMs and employs a separate model for text-to-audio (T2A) generation, which inevitably loses critical visual details and motion dynamics essential for realistic sound synthesis. More recently, DeepSound-V1 (Liang et al., 2025), while introducing MLLM-generated CoT for V2A, fragments the process into three separate tasks (audio generation, vocal removal, and silence detection) using specialized models rather than a unified framework. This fragmentation fails to fully leverage the deep understanding and reasoning capabilities that MLLMs could bring to comprehensive audio design.
@@
  • A unified multimodal foundation model capable of high-quality audio synthesis from arbitrary combination of video, text, and audio inputs, leveraging CoT instructions to decompose complex scenes into manageable sound components.
@@ 2.1 Video-to-Audio Generation
V2A  (Cheng et al., 2024a; Wang et al., 2024c; Xu et al., 2024; Liu et al., 2025) focuses on synthesizing audio that aligns seamlessly with the visual content of a video clip. Earlier work (Luo et al., 2024; Xing et al., 2024; Zhang et al., 2024; Viertola et al., 2024) try to generate audio samples based on silent video only using latent diffusion models (Rombach et al., 2022) or language models (Floridi & Chiriatti, 2020). For example, Diff-Foley (Luo et al., 2024) employs an audio-visual contrastive feature and latent diffusion to predict the spectrogram latent, while FoleyGen uses autoregressive techniques. However, recent research (Chen et al., 2024; Mo et al., 2024; You et al., 2025) pay more attention to generating audios using multimodal control including videos, audios, and texts. For example, MMAudio  (Polyak et al., 2024) uses flow matching conditioned on multi-modal inputs, including videos and texts. MultiFoley (Wu et al., 2024) further uses audio context as an additional input. Despite these advancements, existing work struggles to reason beyond simple object recognition and fails to conduct a step-by-step interactive process with users for audio generation. In this work, we propose ThinkSound, a V2A model with CoT reasoning in MLLMs, which supports step-by-step and interactive audio generation and editing.
@@ 2.2 Large Language Models and Reasoning
LLMs (Liu et al., 2024a; Guo et al., 2025; Hurst et al., 2024) have demonstrated remarkable reasoning capabilities through CoT prompting, enabling complex problem decomposition via intermediate reasoning steps. This paradigm, pioneered by (Wei et al., 2022), has been extended to MLLMs (Alayrac et al., 2022; Yang et al., 2023; Chu et al., 2024) that integrate visual, audio, and textual understanding through cross-modal alignment. Recent works (Rubenstein et al., 2023; Li et al., 2023; Wu et al., 2024) have explored MLLMs’ potential in multimodal reasoning, particularly in visual-audio-textual grounding and cross-modal causal reasoning. Despite these, MLLMs remain under-explored for audio generation. While models like SonicVisionLM (Xie et al., 2024) and DeepSound-V1 (Liang et al., 2025) incorporate V2A, they lack mechanisms to decompose user intent into semantic and temporal reasoning steps. Our ThinkSound proposes a novel framework for audio generation and editing that decomposes the complex task into foundation foley generation, interactive object-centric refinement, and targeted audio editing (Wang et al., 2023; Liu et al., 2024d).
@@ 2.3 Flow Matching
Flow Matching has emerged as a powerful alternative to diffusion models for high-quality audio synthesis (Evans et al., 2024; Liu et al., 2024b, 2023a), directly learning continuous normalizing flows between distributions by training a time-dependent vector field. Recent works Lipman et al. (2022); Liu et al. (2022) establish its theoretical foundations, while Liu et al. (2024c) demonstrated its advantages for audio generation with improved quality and faster sampling. Polyak et al. (2024) further advance this approach by applying conditional flow matching to multimodal audio generation, showing its effectiveness in maintaining temporal coherence across complex audio signals. Our work extends these advances through CoT-guided generation and a novel multi-guidance strategy that integrates control signals from multiple modalities. Our modality-agnostic training approach with classifier-free guidance dropout allows the model to generate high-quality audio from flexible input combinations, addressing both control precision and data scarcity challenges that have limited previous approaches.
@@ 3.1 Multimodal Data Sources
The AudioCoT dataset comprises both video-audio and audio-text pairs. For video-audio data, we utilize VGGSound (Chen et al., 2020) and a curated, non-speech subset of AudioSet (Gemmeke et al., 2017) to ensure broad coverage of real-world audiovisual events. For audio-text data, we aggregate pairs from AudioSet-SL (Hershey et al., 2021), Freesound (Fonseca et al., 2017), AudioCaps (Kim et al., 2019), and BBC Sound Effects  [^1] https://sound-effects.bbcrewind.co.uk/, resulting in a diverse and representative corpus for training multimodal models. Further details on data processing and statistics are provided in the Supplementary Materials’s Section A.1.
@@ 3.2 Automated CoT Generation Pipeline
Stage 1: Foundational Foley CoT Generation. To construct the foley generation CoT data, we employ a multi-stage processing pipeline. For video-audio pairs, we utilize VideoLLaMA2 (Cheng et al., 2024b) to extract both temporal and semantic information from the videos through different prompting strategies, while for audio-text pairs (which lack video data), we employ a more streamlined approach without VideoLLaMA2. In both cases, we generate audio descriptions using Qwen2-Audio (Chu et al., 2024), combining them with existing video captions or text annotations. The collected information is then integrated using GPT-4.1-nano [^2] https://openai.com/index/gpt-4-1/ to synthesize comprehensive CoT reasoning chains that capture the complex relationships between content conditions and the corresponding audio elements, ensuring both data types contribute to a comprehensive understanding of audio generation reasoning.
Stage 2: Interactive Object-Centric CoT Generation. To facilitate object-focused audio generation, we develop a Region of Interest (ROI) extraction framework leveraging Grounded SAM2 (Ren et al., 2024; Ravi et al., 2024). Using audio captions as prompts, we identify and generate bounding boxes for potential sound-emitting objects. These coordinates are tracked temporally across video frames, while VideoLLaMA2 provides detailed semantic descriptions for each ROI segment. For complex audio manipulations (extraction/removal), we construct a hierarchical reasoning structure where the CoT of the target video is merged with another video’s CoT to establish a global context. This composite representation is then integrated with ROI-specific generation information and processed by GPT-4.1-nano to formulate coherent manipulation rationales (detailed in Supplementary Materials A.2).
@@ 4.1 Overview
Figure 3: Overview of the ThinkSound architecture. Left: our Multimodal LLM framework, where a fine-tuned VideoLLaMA 2 model generates CoT reasoning for audio generation and editing. Right: our enhanced Multimodal Transformer architecture, which features an MM-DiT backbone with dedicated pathways for processing multimodal inputs and CoT-driven conditioning to enable high-fidelity, contextually grounded audio generation.
@@ 4.2 CoT Reasoning with Multimodal LLM
To enable stepwise, context-aware audio generation, we leverage VideoLLaMA2 (Cheng et al., 2024b) as the core multimodal reasoning engine. VideoLLaMA2 is selected for its state-of-the-art capability in fusing video, text, and audio modalities, and its advanced spatial-temporal modelling is essential for capturing the nuanced interplay between visual events and their corresponding acoustic manifestations.
@@ 4.3 CoT-Guided Unified Audio Foundation Model
The core of ThinkSound is our unified audio foundation model that seamlessly translates CoT reasoning into high-quality audio, whose detail is shown in the right part of Fig. 3. We encode audio into latent representations using a pre-trained VAE (Pinheiro Cinelli et al., 2021) and train our model with conditional flow matching (Lipman et al., 2022), where the velocity field prediction is conditioned on multimodal context including visual content, CoT reasoning, text captions, and audio context. To support any combination of input modalities, we adopt the integration of classifier-free guidance dropout during training. By randomly dropping different modality combinations with probability $p_{\mathrm{drop}}$, we enable the model to handle arbitrary input configurations at inference time—essential for our interactive framework. We also incorporate strategic audio context masking to support advanced editing operations such as inpainting and extension.
For text processing, we employ a dual-pathway encoding strategy: MetaCLIP (Xu et al., 2023) encodes visual captions to provide scene-level context, while T5-v1-xl (Raffel et al., 2020) processes structured CoT reasoning to capture detailed temporal and causal relationships. These complementary representations are effectively combined, with MetaCLIP’s features serving as global conditioning signals while T5’s outputs enable fine-grained reasoning-driven control.
Our enhanced MM-DiT architecture builds on recent advances in multimodal generative modeling (Esser et al., 2024; Labs, 2024; Cheng et al., 2024a) with three key components: (1) we implement a hybrid transformer backbone that alternates between modality-specific and shared processing. Multi-stream transformer blocks maintain separate parameters for each modality while sharing attention mechanisms, allowing efficient processing of diverse inputs without sacrificing cross-modal learning. (2) we design an adaptive fusion module that upsamples video features and fuses them with audio latents via a gated mechanism (Cho et al., 2014). This not only highlights salient visual cues and suppresses irrelevant information, but also ensures that video information is directly involved in subsequent single-stream transformer blocks. By integrating video into the audio latent space, the model can better capture subtle visual details and their nuanced effects on the soundscape, enabling richer cross-modal reasoning than using audio latents alone. (3) we implement global conditioning by mean-pooling CLIP features from both the caption and video, and, following MMAudio (Cheng et al., 2024a), incorporating sync features to improve audio-visual temporal alignment. The resulting global condition is added to the timestep embedding and injected via adaptive layer normalization layers (AdaLN) (Peebles & Xie, 2023) into both multi-stream and single-stream blocks.
@@ 4.4 Step-by-Step CoT-Guided Audio Generation and Editing
By enabling flexible combinations of input modalities along with CoT, ThinkSound supports decomposing audio generation into three intuitive stages shown in Fig. 1. This three-stage pipeline enables progressively refined, highly customized audio generation through an intuitive interactive workflow, with CoT reasoning bridging user intent and audio synthesis at each step.
@@
In the final stage, users can provide high-level editing instructions to refine audio quality or modify specific elements. The MLLM translates these natural language instructions into precise audio processing operations through CoT reasoning, considering both visual content and current audio state. The foundation model, conditioned on both this reasoning and the existing audio context, applies targeted modifications while maintaining overall coherence. This natural language-driven approach enables non-technical users to perform sophisticated audio manipulation including add sounds, remove sound, audio inpainting, and audio extension.

### v3

@@ Abstract
While end-to-end video-to-audio generation has greatly improved, producing high-fidelity audio that authentically captures the nuances of visual content remains challenging. Like professionals in the creative industries, this generation requires sophisticated reasoning about items such as visual dynamics, acoustic environments, and temporal relationships. We present ThinkSound, a novel framework that leverages Chain-of-Thought (CoT) reasoning to enable stepwise, interactive audio generation and editing for videos. Our approach decomposes the process into three complementary stages: foundational foley generation that creates semantically coherent soundscapes, interactive object-centric refinement through precise user interactions, and targeted editing guided by natural language instructions. At each stage, a multimodal large language model generates contextually aligned CoT reasoning that guides a unified audio foundation model. Furthermore, we introduce AudioCoT, a comprehensive dataset with structured reasoning annotations that establishes connections between visual content, textual descriptions, and sound synthesis. Experiments demonstrate that ThinkSound achieves state-of-the-art performance in video-to-audio generation across both audio metrics and CoT metrics, and excels in the out-of-distribution Movie Gen Audio benchmark. The project page is available at https://ThinkSound-Project.github.io.
@@ 1 Introduction
Figure 1: ThinkSound with CoT: (1) CoT-driven foley synthesis captures semantic and temporal details (2) interactive object-centric refinement for control (3) targeted editing.
Generating realistic sound for video demands more than recognizing objects; it requires reasoning about complex visual dynamics and context – determining when an owl is chirping versus flapping its wings, identifying the subtle sway of tree branches, and synchronizing multiple sound events within a scene, as illustrated in Figure 1. Current end-to-end video-to-audio (V2A) generation systems (luo2024difffoley; xing2024seeing; zhang2024foleycrafter), while having largely improved, often struggle with this compositional complexity and contextual nuance. They may produce generic sounds or fail to synchronize precisely with subtle visual cues, thus limiting fidelity and user control.
Recent breakthroughs in Multimodal Large Language Models (MLLMs) (lu2024deepseek; liu2023visual; wang2024qwen2) offer powerful capabilities in understanding and reasoning across multiple modalities. Concurrently, Chain-of-Thought (CoT) prompting (wei2022chain; zhang2023multimodal; wang2025multimodal) has proven effective at eliciting structured, step-by-step reasoning from large models. These advancements create great potential to fundamentally rethink V2A by decomposing complex sound design into explicit reasoning steps and actionable synthesis instructions. The necessity for CoT in V2A becomes evident when examining the process of sound designers. They employ a multi-stage approach: analyzing visual content, then reasoning about acoustic properties, and finally synthesizing and refining sounds. End-to-end approaches compress the process into a single black-box transformation, losing the nuanced reasoning required for audio generation.
Some approaches have attempted to adopt MLLMs for multi-stage V2A. SonicVisionLM (xie2024sonicvisionlm) converts video to textual captions using MLLMs and employs a separate model for text-to-audio (T2A) generation, which inevitably loses critical visual details and motion dynamics essential for realistic sound synthesis. More recently, DeepSound-V1 (liang2025deepsoundv1startthinkstepbystep), while introducing MLLM-generated CoT for V2A, fragments the process into three separate tasks (audio generation, vocal removal, and silence detection) using specialized models rather than a unified framework. This fragmentation fails to fully leverage the deep understanding and reasoning capabilities that MLLMs could bring to comprehensive audio design.
@@
  • A unified multimodal foundation model capable of high-quality audio synthesis from an arbitrary combination of video, text, and audio inputs, leveraging CoT instructions to decompose complex scenes into manageable sound components.
@@ 2.1 Video-to-Audio Generation
V2A  (cheng2024taming; wang2024tiva; xu2024video; liu2025omniaudio) focuses on synthesizing audio that aligns seamlessly with the visual content of a video clip. Earlier work (luo2024difffoley; xing2024seeing; zhang2024foleycrafter; viertola2024temporally) try to generate audio samples based on silent video only using latent diffusion models (rombach2022high) or language models (floridi2020gpt). For example, Diff-Foley (luo2024difffoley) employs an audio-visual contrastive feature and latent diffusion to predict the spectrogram latent, while FoleyGen uses autoregressive techniques. However, recent research (chen2024video; mo2024text; you2025ta) pay more attention to generating audios using multimodal control, including videos, audios, and texts. For example, MMAudio  (polyak2024movie) uses flow matching conditioned on multi-modal inputs, including videos and texts. MultiFoley (wu2024next) further uses audio context as an additional input. Despite these advancements, existing work struggles to reason beyond simple object recognition and fails to conduct a step-by-step interactive process with users for audio generation. In this work, we propose ThinkSound, a V2A model with CoT reasoning in MLLMs, which supports step-by-step and interactive audio generation and editing.
@@ 2.2 Large Language Models and Reasoning
LLMs (liu2024deepseek; guo2025deepseek; hurst2024gpt) have demonstrated remarkable reasoning capabilities through CoT prompting, enabling complex problem decomposition via intermediate reasoning steps. This paradigm, pioneered by (wei2022chain), has been extended to MLLMs (alayrac2022flamingo; yang2023dawn; chu2024qwen2) that integrate visual, audio, and textual understanding through cross-modal alignment. Recent works (rubenstein2023audiopalm; li2023videochat; wu2024next) have explored MLLMs’ potential in multimodal reasoning, particularly in visual-audio-textual grounding and cross-modal causal reasoning. Despite these, MLLMs remain under-explored for audio generation. While models like SonicVisionLM (xie2024sonicvisionlm) and DeepSound-V1 (liang2025deepsoundv1startthinkstepbystep) incorporate V2A, they lack mechanisms to decompose user intent into semantic and temporal reasoning steps. Our ThinkSound proposes a novel framework for audio generation and editing that decomposes the complex task into foundation Foley generation, interactive object-centric refinement, and targeted audio editing (wang2023audit; medicliu).
@@ 2.3 Flow Matching
Flow Matching has emerged as a powerful alternative to diffusion models for high-quality audio synthesis (evans2024fast; liu2024audiolcm; liu2023vit), directly learning continuous normalizing flows between distributions by training a time-dependent vector field. Recent works lipman2022flow; liu2022flow establish its theoretical foundations, while liu2024flashaudio demonstrated its advantages for audio generation with improved quality and faster sampling. polyak2024movie further advances this approach by applying conditional flow matching to multimodal audio generation, showing its effectiveness in maintaining temporal coherence across complex audio signals. Our work extends these advances through CoT-guided generation and a novel multi-guidance strategy that integrates control signals from multiple modalities. Our modality-agnostic training approach with classifier-free guidance dropout allows the model to generate high-quality audio from flexible input combinations, addressing both control precision and data scarcity challenges that have limited previous approaches.
@@ 3.1 Multimodal Data Sources
The AudioCoT dataset comprises both video-audio and audio-text pairs. For video-audio data, we utilize VGGSound (chen2020vgg) and a curated, non-speech subset of AudioSet (gemmeke2017audio) to ensure broad coverage of real-world audiovisual events. For audio-text data, we aggregate pairs from AudioSet-SL (hershey2021benefit), Freesound (fonseca2017freesound), AudioCaps (kim2019audiocaps), and BBC Sound Effects  [^1] https://sound-effects.bbcrewind.co.uk/, resulting in a diverse and representative corpus for training multimodal models. Further details on data processing and statistics are provided in the Appendix A.1.
@@ 3.2 Automated CoT Generation Pipeline
Stage 1: Foundational Foley CoT Generation. To construct the Foley generation CoT data, we employ a multi-stage processing pipeline. For video-audio pairs, we utilize VideoLLaMA2 (cheng2024videollama) to extract both temporal and semantic information from the videos through different prompting strategies, while for audio-text pairs (which lack video data), we employ a more streamlined approach without VideoLLaMA 2. In both cases, we generate audio descriptions using Qwen2-Audio (chu2024qwen2), combining them with existing video captions or text annotations. The collected information is then integrated using GPT-4.1-nano [^2] https://openai.com/index/gpt-4-1/ to synthesize comprehensive CoT reasoning chains that capture the complex relationships between content conditions and the corresponding audio elements, ensuring both data types contribute to a comprehensive understanding of audio generation reasoning.
Stage 2: Interactive Object-Centric CoT Generation. To facilitate object-focused audio generation, we develop a Region of Interest (ROI) extraction framework leveraging Grounded SAM2 (ren2024grounding; ravi2024sam2segmentimages). Using audio captions as prompts, we identify and generate bounding boxes for potential sound-emitting objects. These coordinates are tracked temporally across video frames, while VideoLLaMA2 provides detailed semantic descriptions for each ROI segment. For complex audio manipulations (extraction/removal), we construct a hierarchical reasoning structure where the CoT of the target video is merged with another video’s CoT to establish a global context. This composite representation is then integrated with ROI-specific generation information and processed by GPT-4.1-nano to formulate coherent manipulation rationales (detailed in Appendix A.2).
@@ 4.1 Overview
Figure 3: Overview of the ThinkSound architecture. Left: our Multimodal LLM framework, where a fine-tuned VideoLLaMA 2 model generates CoT reasoning for audio generation and editing. Right: our enhanced Multimodal Transformer architecture, which features an $M$-DiT backbone with dedicated pathways for processing multimodal inputs and CoT-driven conditioning to enable high-fidelity, contextually grounded audio generation.
@@ 4.2 CoT Reasoning with Multimodal LLM
To enable stepwise, context-aware audio generation, we leverage VideoLLaMA2 (cheng2024videollama) as the core multimodal reasoning engine. VideoLLaMA2 is selected for its state-of-the-art capability in fusing video, text, and audio modalities, and its advanced spatial-temporal modelling is essential for capturing the nuanced interplay between visual events and their corresponding acoustic manifestations.
@@ 4.3 CoT-Guided Unified Audio Foundation Model
The core of ThinkSound is our unified audio foundation model that seamlessly translates CoT reasoning into high-quality audio, whose detail is shown in the right part of Figure 3. We encode audio into latent representations using a pre-trained VAE (pinheiro2021variational) and train our model with conditional flow matching (lipman2022flow), where the velocity field prediction is conditioned on multimodal context, including visual content, CoT reasoning, text captions, and audio context. To support any combination of input modalities, we adopt the integration of classifier-free guidance dropout during training. By randomly dropping different modality combinations with probability $p_{\mathrm{drop}}$, we enable the model to handle arbitrary input configurations at inference time—essential for our interactive framework. We also incorporate strategic audio context masking to support advanced editing operations such as inpainting and extension.
For text processing, we employ a dual-pathway encoding strategy: MetaCLIP (xu2023demystifying) encodes visual captions to provide scene-level context, while T5-v1-xl (2020t5) processes structured CoT reasoning to capture detailed temporal and causal relationships. These complementary representations are effectively combined, with MetaCLIP’s features serving as global conditioning signals while T5’s outputs enable fine-grained reasoning-driven control.
Our enhanced MM-DiT architecture builds on recent advances in multimodal generative modeling (esser2024scaling; flux2024; cheng2024taming) with three key components: (1) we implement a hybrid transformer backbone that alternates between modality-specific and shared processing. Multi-stream transformer blocks maintain separate parameters for each modality while sharing attention mechanisms, allowing efficient processing of diverse inputs without sacrificing cross-modal learning. (2) We design an adaptive fusion module that upsamples video features and fuses them with audio latents via a gated mechanism (cho2014learning). This not only highlights salient visual cues and suppresses irrelevant information, but also ensures that video information is directly involved in subsequent single-stream transformer blocks. By integrating video into the audio latent space, the model can better capture subtle visual details and their nuanced effects on the soundscape, enabling richer cross-modal reasoning than using audio latents alone. (3) We implement global conditioning by mean-pooling CLIP features from both the caption and video, and, following MMAudio (cheng2024taming), incorporating sync features to improve audio-visual temporal alignment. The resulting global condition is added to the timestep embedding and injected via adaptive layer normalization layers (AdaLN) (peebles2023scalable) into both multi-stream and single-stream blocks.
@@ 4.4 Step-by-Step CoT-Guided Audio Generation and Editing
By enabling flexible combinations of input modalities along with CoT, ThinkSound supports decomposing audio generation into three intuitive stages shown in Figure 1. This three-stage pipeline enables progressively refined, highly customized audio generation through an intuitive interactive workflow, with CoT reasoning bridging user intent and audio synthesis at each step.
@@
In the final stage, users can provide high-level editing instructions to refine audio quality or modify specific elements. The MLLM translates these natural language instructions into precise audio processing operations through CoT reasoning, considering both visual content and current audio state. The foundation model, conditioned on both this reasoning and the existing audio context, applies targeted modifications while maintaining overall coherence. This natural language-driven approach enables non-technical users to perform sophisticated audio manipulation, including adding sounds, removing sounds, audio inpainting, and audio extension.